# Dynamic Modes as Time Representation for Spatiotemporal Forecasting

### v1

@@ Abstract
1 Introduction
Spatiotemporal forecasting plays a central role in diverse real-world applications, including traffic flow prediction (Vlahogianni et al., 2014), climate modeling (Reichstein et al., 2019), energy demand estimation (Hong et al., 2016), and environmental monitoring (Pettorelli et al., 2014). These systems often exhibit strong periodic or seasonal patterns—such as daily commuting cycles, weekly temperature fluctuations, or yearly rainfall changes—that are essential for accurate long-term forecasting and planning. However, capturing such long-range periodic structures remains a persistent challenge for spatiotemporal forecasting (Vlahogianni et al., 2014; Karlaftis and Vlahogianni, 2011).
Recent advances in deep learning have enabled powerful spatiotemporal forecasting models by combining neural sequence architectures and spatial representations, such as Recurrent Neural Networks (RNNs) (Hochreiter and Schmidhuber, 1997; Chung et al., 2014), Temporal Convolutional Networks (TCNs) (Yu and Koltun, 2016), Graph Convolutional Networks (GCNs) (Defferrard et al., 2016; Kipf and Welling, 2016), and Transformer-based models (Cai et al., 2020; Grigsby et al., 2021). While these models effectively capture short-term dynamics, they often struggle to represent global temporal dependencies unless aided by explicit temporal encodings. A common workaround is to include hand-crafted time features (e.g., time-of-day, day-of-week(Kazemi et al., 2019), but such features are often domain-specific and insufficient to express complex multi-scale seasonality.
To address this limitation, we propose a data-driven approach to extract and encode periodic structure directly from observations, using Dynamic Mode Decomposition (DMD) (Schmid, 2022). By leveraging Koopman operator theory, DMD decomposes spatiotemporal signals into interpretable oscillatory modes without requiring timestamp metadata. We construct time embeddings by extracting the real and imaginary components of the dominant DMD modes (Brunton et al., 2016a; Jovanović et al., 2014), resulting in a compact spectral representation of temporal dynamics. These embeddings can be seamlessly integrated as covariates into any forecasting model to enhance its ability to capture global periodic dependencies.
@@ 1 Introduction
We evaluate our method on three real-world spatiotemporal datasets: GZ-METRO (urban metro ridership) (Li et al., 2018), PEMS04 (highway traffic flow) (Song et al., 2020), and Daymet (daily climate temperature) (Thornton et al., 2022). We integrate our DMD-based time embeddings into several deep learning models and demonstrate improved long-term forecasting accuracy, reduced residual correlations, and enhanced temporal generalization across diverse settings.
The rest of the paper is organized as follows. Section 2 reviews related work on spatiotemporal forecasting and time representations. Section 3 presents our proposed method. Section 4 describes the experiments and analysis. Section 5 concludes the paper.
2 Related Work
2.1 Spatiotemporal Forecasting Models
@@
Statistical and linear algebraic models. Classical statistical models analyze time series in either the time domain (e.g., ARMA, ARIMA, SARIMA (Box et al., 2015)) or frequency domain (e.g., Fourier and wavelet transforms (Cryer and Chan, 2008)). Matrix factorization techniques have also been applied to low-rank spatiotemporal forecasting and imputation tasks (Yu et al., 2016).
Dynamic systems models. DMD (Schmid, 2022; Avila and Mezić, 2020) and Sparse Identification of Nonlinear Dynamics (SINDy) (Brunton et al., 2016b; Champion et al., 2019a) aim to recover governing structures from observations using data-driven operator-based formulations.
Machine learning models. Early approaches include tree ensembles and support vector machines (Bishop and Nasrabadi, 2006), while deep learning (DL) models have become the dominant paradigm for capturing complex spatiotemporal dependencies (LeCun et al., 2015). Our work contributes to this direction by enhancing DL models with a novel spectral time embedding.
2.2 Deep Learning for Spatiotemporal Forecasting
DL-based spatiotemporal models combine temporal encoders and spatial graph modules. DCRNN (Li et al., 2018) integrates diffusion convolution into GRUs to model both space and time. STGCN (Yu et al., 2018) uses GCNs for spatial structure and CNNs for temporal encoding. ASTGCN (Guo et al., 2019) adds spatial-temporal attention to the STGCN backbone.
Graph WaveNet (Wu et al., 2019) introduces an adaptive adjacency matrix and dilated temporal convolutions, offering flexibility in learning the spatial graph structure. STSGCN (Song et al., 2020) models temporal adjacency by linking graphs across consecutive time steps. GMAN (Zheng et al., 2020) adopts dual attention over space and time with embedding fusion. FC-GAGA (Oreshkin et al., 2021) extends N-BEATS (Oreshkin et al., 2020) with a gated graph layer that learns sparse, non-Markovian structures.
@@ 2.2 Deep Learning for Spatiotemporal Forecasting
2.3 Time Embedding and Periodic Structure Learning
Encoding long-term temporal dependency is critical in spatiotemporal forecasting due to prevalent seasonal patterns. The most common practice is to use hand-crafted features such as time-of-day or day-of-week, which capture fixed daily or weekly periodicities. However, these features are rigid and may not reflect latent periodic structures in the data.
Time2Vec (Kazemi et al., 2019) extends sinusoidal positional encoding (Vaswani et al., 2017) by learning frequencies and phase shifts, offering a flexible way to encode global time. This method has been widely adopted in forecasting tasks (Grigsby et al., 2021) and is related to Fourier feature mappings (Tancik et al., 2020; Li et al., 2021). However, recent studies suggest that Fourier features often overfit high-frequency fluctuations and struggle to encode low-frequency seasonality (Tancik et al., 2020).
In contrast, we propose a spectral embedding approach based on DMD. Our method extracts dominant oscillatory modes from observed time series via Koopman operator analysis (Brunton et al., 2016a; Jovanović et al., 2014). The real and imaginary components of these modes form a compact and interpretable time representation, which captures complex and data-driven periodic patterns beyond hand-crafted or fixed-frequency alternatives. These embeddings can be seamlessly integrated into any DL forecasting model to enhance long-term temporal modeling.
3 Methodology
@@ 3.1 Spatiotemporal Forecasting
3.2 Spectral Analysis of Spatiotemporal Data with DMD
@@
    Hankel embedding to expand the effective spatial dimension of the data (since typically $N\ll T$ in spatiotemporal forecasting) (Champion et al., 2019b; Kamb et al., 2020; Wang and Sun, 2022);
@@
    Total DMD (TDMD) for unbiased, noise-aware spectral decomposition (Hemati et al., 2017);
@@
    Sparsity-Promoting DMD (SPDMD) to select dominant dynamic modes (Jovanović et al., 2014).
Our goal is to obtain a compact and interpretable set of oscillatory components whose temporal evolution will serve as covariates in downstream forecasting models.
@@
where $f$ is an unknown nonlinear function. To linearize the system, we follow the Koopman operator framework (Schmid, 2022), which lifts the dynamics into a higher-dimensional observable space via a function $\phi(\cdot)$ such that:
@@
To encourage interpretability, we apply SPDMD (Jovanović et al., 2014) to prune to the $r$ dominant modes, yielding:
@@
To improve the performance of DMD on high-dimensional data, we first enhance the data using a circulant Hankel embedding (Schmid, 2022; Wang and Sun, 2023), which converts $\mathbf{D}$ into a tall matrix:
@@
Since directly computing the SVD of $\mathbf{H}\mathbf{H}^{\top}\in\mathbb{R}^{N\tau\times N\tau}$ is computationally expensive for large $\tau$, we apply the method of snapshots (Brunton and Kutz, 2022), which leverages the symmetry of the inner product space to reduce the computation to a smaller Gram matrix. Specifically, we perform the truncated SVD on the smaller matrix $\mathbf{H}^{\top}\mathbf{H}\in\mathbb{R}^{T\times T}$:
@@
where $\tilde{\mathbf{V}}\in\mathbb{R}^{T\times r}$ contains the top-$r$ right singular vectors; $\tilde{\mathbf{\Sigma}}\in\mathbb{R}^{r\times r}$ is the diagonal matrix of singular values.
@@
which approximates the linear transition matrix in the low-rank temporal subspace. The eigenvalues $\{\lambda_{1},\dots,\lambda_{r}\}$ of $\mathbf{A}_{\text{DMD}}$ characterize the oscillatory behavior of the system (magnitude $\rightarrow$ growth/decay, imaginary part $\rightarrow$ frequency) (Schmid, 2022).
Figure 1: Visualization of temporal modes extracted via DMD for the GZ-METRO dataset. Each mode evolves as a complex exponential $e^{i2\pi ft}$, with real (cosine), imaginary (sine), and modulus components. Left: high-frequency mode; Right: low-frequency mode. The real and imaginary parts form the final time embedding.
@@
Each row in $\bar{\mathbf{C}}$ defines the temporal profile of a dominant DMD mode. These complex-valued signals form the basis of our time embedding, as illustrated in Figure 1.
@@
At each time step $t$, we extract the real and imaginary parts of $\bar{\mathbf{C}}[:,t]$ to construct a $2r$-dimensional time covariate:
@@
4 Experiments
To evaluate the effectiveness of the proposed method, we carried out experiments on three spatiotemporal datasets: GZ-METRO, PEMS04, and Daymet. GZ-METRO records 15-min ridership based on smart card tap-in data at 159 metro stations for 3 months in Guangzhou, China. PEMS04 is a highway traffic flow datasets used in (Song et al., 2020). The dataset records 5-min aggregated traffic flow from the Cal-trans Performance Measurement System (PeMS) (Chen et al., 2001). Daymet is a dataset that provides the daily maximum temperature recordings for Continental North America from 2010 to 2021 (Thornton et al., 2022). The summary of datasets can be found in Table 1. For GZ-METRO and Daymet, 70% of the data were used for training, 20% for testing, and the remaining 10% for validation. For PEMS04, the split ratio is 6:2:2 as used in the original paper (Song et al., 2020). All datasets were applied z-score normalization with statistics obtained from the training set. We selected several spatiotemporal forecasting frameworks as the base models:
  • FC-LSTM (Sutskever et al., 2014): A Sequence-to-Sequence LSTM model with fully-connected LSTM layers in both encoder and decoder.
  • DCRNN (Li et al., 2018): Diffusion convolution recurrent neural network, which wraps diffusion convolution operation into recurrent neural networks to achieve spatiotemporal forecasting.
  • Graph WaveNet (Wu et al., 2019): A spatiotemporal forecasting model that combines dilated 1D convolution for modeling temporal dynamics and graph convolution for modeling spatial dynamics.
  • AGCRN (Bai et al., 2020): Similar to Graph WaveNet which uses an adaptive adjacency matrix, the model uses GRU to model temporal dependency.
@@ 4.1 Quantitative Study
4.2 Qualitative Study
4.2.1 Effectiveness of DMD-based time embedding
To evaluate the effectiveness of our DMD-based time embedding, we compare it with several alternatives for encoding periodic patterns: (i) time of day (D), (ii) time of day, day of week (DW), and (iii) the learnable sinusoidal embedding method Time2Vec (Kazemi et al., 2019).
Our observations align with the findings of (Tancik et al., 2020), which showed that learned Fourier-based time features tend to concentrate on high-frequency components and struggle to capture long-range periodicity. We further inspected the learned frequencies of Time2Vec and found that they mostly correspond to short-term fluctuations with periods between 5 and 15 time steps (i.e., 25–75 minutes), indicating a failure to model daily or weekly cycles. By contrast, our DMD-based embedding captures multiple frequencies of varying scales, including long-term modes, as visualized in Figure 1.
4.2.2 Residual correlation and temporal dependency analysis
Figure 3(a) and Figure 3(b) show the residual correlation matrices computed from the 12-step-ahead prediction errors at three time lags: $S=0$, $72$, and $504$ steps, corresponding to contemporaneous, daily, and weekly intervals. Each panel visualizes the correlations between residual vectors $\bm{\eta}_{t}$ and $\bm{\eta}_{t-S}$ over all spatiotemporal positions. Without our method, we observe strong overall correlations at $S=0$ (0.138), and noticeable dependencies at $S=72$ (0.081) and $S=504$ (0.096). In contrast, our method reduces these values to 0.104, 0.063, and 0.079, respectively. These reductions indicate that our model effectively captures structured periodic patterns and reduces error autocorrelation across periods.
To further validate this observation, we analyze the autocorrelation function (ACF) of the prediction residuals from three representative stations (A, B, and C). Figure 4(a) shows the ACF plots of the original Graph WaveNet model. We observe clear peaks around lag 72 and 144, corresponding to daily and bi-daily periodicity, suggesting that the model fails to fully capture these regular patterns, which manifest as cyclic structure in the residuals. In comparison, Figure 4(b) illustrates the ACF of the residuals using Graph WaveNet augmented with our time embedding. The periodic peaks at lag 72 and 144 are substantially diminished across all three stations. This indicates that the DMD-based time features help the model internalize the dominant periodic components in the data, leading to less temporally correlated residuals and improved forecasting generalization.
@@ 4.2.2 Residual correlation and temporal dependency analysis
Figure 3: Residual correlation matrices of GZ-METRO at time lags $S=0$, $72$, and $504$. Each panel shows absolute residual correlation values between spatiotemporal locations. Values in parentheses indicate mean absolute correlation.
(a) Without time embedding (Graph WaveNet only).
(b) With our method (Graph WaveNet + DMD).
Figure 4: Autocorrelation functions (ACF) of 12-step prediction residuals from three representative sensors in GZ-METRO. Peaks around lag 72 indicate periodic errors; these are reduced by our time embedding.
5 Conclusion

### v2

@@ Abstract
Introduction
Spatiotemporal forecasting plays a central role in diverse real-world applications, including traffic flow prediction (Vlahogianni, Karlaftis, and Golias 2014), climate modeling (Reichstein et al. 2019), energy demand estimation (Hong, Pinson, and Fan 2016), and environmental monitoring (Pettorelli et al. 2014). These systems often exhibit strong periodic or seasonal patterns—such as daily commuting cycles, weekly temperature fluctuations, or yearly rainfall changes—that are essential for accurate long-term forecasting and planning. However, capturing such long-range periodic structures remains a persistent challenge for spatiotemporal forecasting (Vlahogianni, Karlaftis, and Golias 2014; Karlaftis and Vlahogianni 2011).
Recent advances in deep learning have enabled powerful spatiotemporal forecasting models by combining neural sequence architectures and spatial representations, such as Recurrent Neural Networks (RNNs) (Hochreiter and Schmidhuber 1997; Chung et al. 2014), Temporal Convolutional Networks (TCNs) (Yu and Koltun 2016), Graph Convolutional Networks (GCNs) (Defferrard, Bresson, and Vandergheynst 2016; Kipf and Welling 2016), and Transformer-based models (Cai et al. 2020; Grigsby, Wang, and Qi 2021). While these models effectively capture short-term dynamics, they often struggle to represent global temporal dependencies unless aided by explicit temporal encodings. A common workaround is to include hand-crafted time features (e.g., time-of-day, day-of-week(Kazemi et al. 2019), but such features are often domain-specific and insufficient to express complex multi-scale seasonality.
To address this limitation, we propose a data-driven approach to extract and encode periodic structure directly from observations, using Dynamic Mode Decomposition (DMD) (Schmid 2022). By leveraging Koopman operator theory, DMD decomposes spatiotemporal signals into interpretable oscillatory modes without requiring timestamp metadata. We construct time embeddings by extracting the real and imaginary components of the dominant DMD modes (Brunton et al. 2016; Jovanović, Schmid, and Nichols 2014), resulting in a compact spectral representation of temporal dynamics. These embeddings can be seamlessly integrated as covariates into any forecasting model to enhance its ability to capture global periodic dependencies.
@@ Introduction
We evaluate our method on three real-world spatiotemporal datasets: GZ-METRO (urban metro ridership) (Li et al. 2018), PEMS04 (highway traffic flow) (Song et al. 2020), and Daymet (daily climate temperature) (Thornton et al. 2022). We integrate our DMD-based time embeddings into several deep learning models and demonstrate improved long-term forecasting accuracy, reduced residual correlations, and enhanced temporal generalization across diverse settings.
Table 1: Comparison of Time Embedding Methods
Method Expressiveness Interpretability Generalization
Time2Vec Learns both periodic and non‐periodic patterns Medium: internal weights not directly interpretable Strong: end‐to‐end learning, no manual design
Fourier Features Fixed‐frequency sinusoids High: clear spectral interpretation Limited: requires predefined frequencies; less robust to non‐stationarity
Handcrafted Features Predefined calendars and events High: semantically intuitive Limited: domain‐specific; poor cross‐task transfer
DMD-based Embedding Automatically extracts multiple periodic modes High: modes correspond to spectral components Strong: fully data‐driven, domain‐agnostic
Related Work
Spatiotemporal Forecasting Models
@@
Statistical and linear algebraic models. Classical statistical models analyze time series in either the time domain (e.g., ARMA, ARIMA, SARIMA (Box et al. 2015)) or frequency domain (e.g., Fourier and wavelet transforms (Cryer and Chan 2008)). Matrix factorization techniques have also been applied to low-rank spatiotemporal forecasting and imputation tasks (Yu, Rao, and Dhillon 2016).
Dynamic systems models. DMD (Schmid 2022; Avila and Mezić 2020) and Sparse Identification of Nonlinear Dynamics (SINDy) (Brunton, Proctor, and Kutz 2016; Champion et al. 2019) aim to recover governing structures from observations using data-driven operator-based formulations.
Machine learning models. Early approaches include tree ensembles and support vector machines (Bishop and Nasrabadi 2006), while deep learning (DL) models have become the dominant paradigm for capturing complex spatiotemporal dependencies (LeCun, Bengio, and Hinton 2015). Our work contributes to this direction by enhancing DL models with a novel spectral time embedding.
Deep Learning for Spatiotemporal Forecasting
DL-based spatiotemporal models combine temporal encoders and spatial graph modules. DCRNN (Li et al. 2018) integrates diffusion convolution into GRUs to model both space and time. STGCN (Yu, Yin, and Zhu 2018) uses GCNs for spatial structure and CNNs for temporal encoding. ASTGCN (Guo et al. 2019) adds spatial-temporal attention to the STGCN backbone.
Graph WaveNet (Wu et al. 2019) introduces an adaptive adjacency matrix and dilated temporal convolutions, offering flexibility in learning the spatial graph structure. STSGCN (Song et al. 2020) models temporal adjacency by linking graphs across consecutive time steps. GMAN (Zheng et al. 2020) adopts dual attention over space and time with embedding fusion. FC-GAGA (Oreshkin et al. 2021) extends N-BEATS (Oreshkin et al. 2020) with a gated graph layer that learns sparse, non-Markovian structures.
@@
Time Embedding and Periodic Structure Learning
Encoding long-term temporal dependency is critical in spatiotemporal forecasting due to prevalent seasonal patterns. The most common practice is to use hand-crafted features such as time-of-day or day-of-week, which capture fixed daily or weekly periodicities. However, these features are rigid and may not reflect latent periodic structures in the data. Time2Vec (Kazemi et al. 2019) extends sinusoidal positional encoding (Vaswani et al. 2017) by learning frequencies and phase shifts, offering a flexible way to encode global time. This method has been widely adopted in forecasting tasks (Grigsby, Wang, and Qi 2021) and is related to Fourier feature mappings (Tancik et al. 2020; Li et al. 2021). However, recent studies suggest that Fourier features often overfit high-frequency fluctuations and struggle to encode low-frequency seasonality (Tancik et al. 2020).
In contrast, we propose a spectral embedding approach based on DMD. Our method extracts dominant oscillatory modes from observed time series via Koopman operator analysis (Brunton et al. 2016; Jovanović, Schmid, and Nichols 2014). The real and imaginary components of these modes form a compact and interpretable time representation, which captures complex and data-driven periodic patterns beyond hand-crafted or fixed-frequency alternatives. These embeddings can be seamlessly integrated into any DL forecasting model to enhance long-term temporal modeling.
Comparison of Time Embedding Methods.
As Table 1 shows, traditional Fourier and handcrafted time features offer clear interpretability but lack flexibility for complex or non‐stationary signals. Time2Vec can flexibly fit a variety of temporal patterns via end‐to‐end training, yet its internal representations are not readily interpretable. In contrast, our DMD-based embedding automatically identifies and separates dominant spectral modes—providing both high expressiveness and clear modal interpretation—while remaining entirely data‐driven and broadly generalizable.
In the next section, we build on these advantages to introduce our full pipeline for spatiotemporal forecasting.
Figure 1: Proposed pipeline. Raw multivariate signals are first converted into a stacked Hankel matrix and then decomposed via Total TDMD. Sparsity-Promoting SPDMD selects the most prominent eigenmodes; their real and imaginary components yield compact cosine–sine time covariates. These covariates are concatenated with other features and fed into any spatiotemporal forecasting model.
Methodology
@@
Spectral Analysis of Spatiotemporal Data with DMD
@@
    Hankel embedding to expand the effective spatial dimension of the data (since typically $N\ll T$ in spatiotemporal forecasting) (Champion, Brunton, and Kutz 2019; Kamb et al. 2020; Wang and Sun 2022);
@@
    Total DMD (TDMD) for unbiased, noise-aware spectral decomposition (Hemati et al. 2017);
@@
    Sparsity-Promoting DMD (SPDMD) to select dominant dynamic modes (Jovanović, Schmid, and Nichols 2014).
The complete processing chain is summarised in Figure 1. Each stage will be detailed next. Our goal is to obtain a compact and interpretable set of oscillatory components whose temporal evolution will serve as covariates in downstream forecasting models.
@@
where $f$ is an unknown nonlinear function. To linearize the system, we follow the Koopman operator framework (Schmid 2022), which lifts the dynamics into a higher-dimensional observable space via a function $\phi(\cdot)$ such that:
@@
Figure 2 illustrates this idea: a nonlinear spiral trajectory (left) is “lifted” into a high-dimensional space where the Koopman operator acts as a simple rotation; its eigenvalues appear as points on/near the unit circle.
Figure 2: Koopman operator intuition. (A) A nonlinear spiral trajectory is lifted into a higher-dimensional observable space. (B) In that space the dynamics reduce to a simple rotation; the eigenvalues of the Koopman operator lie on (or near) the unit circle, each dot representing an oscillatory mode whose frequency is given by its angle.
To encourage interpretability, we apply SPDMD (Jovanović, Schmid, and Nichols 2014) to prune to the $r$ dominant modes, yielding:
@@
To improve the performance of DMD on high-dimensional data, we first enhance the data using a circulant Hankel embedding (Schmid 2022; Wang and Sun 2023), which converts $\mathbf{D}$ into a tall matrix:
@@
Since directly computing the SVD of $\mathbf{H}\mathbf{H}^{\top}\in\mathbb{R}^{N\tau\times N\tau}$ is computationally expensive for large $\tau$, we apply the method of snapshots (Brunton and Kutz 2022), which leverages the symmetry of the inner product space to reduce the computation to a smaller Gram matrix. Specifically, we perform the truncated SVD on the smaller matrix $\mathbf{H}^{\top}\mathbf{H}\in\mathbb{R}^{T\times T}$:
@@
[…]
@@
which approximates the linear transition matrix in the low-rank temporal subspace. The eigenvalues $\{\lambda_{1},\dots,\lambda_{r}\}$ of $\mathbf{A}_{\text{DMD}}$ characterize the oscillatory behavior of the system (magnitude $\rightarrow$ growth/decay, imaginary part $\rightarrow$ frequency) (Schmid 2022).
@@
Each row in $\bar{\mathbf{C}}$ defines the temporal profile of a dominant DMD mode. These complex-valued signals form the basis of our time embedding.
@@
[…]
@@
Remark.
Each DMD eigenvalue can be written as a complex exponential form $\lambda_{k}=e^{\mu_{k}+i\omega_{k}}$, where $\mu_{k}$ denotes the growth or decay rate, and $\omega_{k}$ denotes the oscillation frequency. In our construction of time covariates, we exclusively use the frequency components $\omega_{k}$ to generate sinusoidal embeddings:
$\mathbf{c}_{t}^{(\text{DMD})}=[\cos(\omega_{1}t),\ldots,\cos(\omega_{K}t),\sin(\omega_{1}t),\ldots,\sin(\omega_{K}t)].$
The amplitude dynamics $e^{\mu_{k}t}$ are discarded to avoid potential numerical instability and to ensure the covariates remain interpretable and stationary. This design decouples trend-like dynamics from periodic components, allowing the main model to focus on learning non-periodic residual structures.
Algorithm 1 DMD-based Time Covariate Construction
0: Raw signals $\{\mathbf{z}_{t}\}_{t=0}^{T-1}$, Hankel window $\tau$, mode count $r$
0: Time covariates $\{\mathbf{c}_{t}^{(\text{DMD})}\}_{t=0}^{T-1}$
1: Build stacked Hankel matrix $\mathbf{H}$ (Eq. 8)
2: Compute SVD: $\mathbf{H}^{\top}\mathbf{H}=\widetilde{\mathbf{V}}\,\widetilde{\mathbf{\Sigma}}^{2}\,\widetilde{\mathbf{V}}^{\top}$ (Eq. 9)
3: Recover $\widetilde{\mathbf{U}}=\mathbf{H}\,\widetilde{\mathbf{V}}\,\widetilde{\mathbf{\Sigma}}^{-1}$ (Eq. 10)
4: Form reduced operator: $\mathbf{A}_{\text{DMD}}=\widetilde{\mathbf{U}}^{\top}\mathbf{H}^{\prime}\,\widetilde{\mathbf{V}}\,\widetilde{\mathbf{\Sigma}}^{-1}$ (Eq. 11)
5: Compute eigenpairs $\{(\lambda_{k},\psi_{k})\}_{k=1}^{r}$ of $\mathbf{A}_{\text{DMD}}$
6:for $t=0$ to $T-1$ do
7:  $\mathbf{c}_{t}^{(\text{DMD})}=[\cos(\omega_{1}t),\dots,\sin(\omega_{r}t)]$ with $\omega_{k}=\arg(\lambda_{k})$
8:end for
9:return  $\{\mathbf{c}_{t}^{(\text{DMD})}\}$
Experiments
To evaluate the effectiveness of the proposed method, we carried out experiments on three spatiotemporal datasets: GZ-METRO, PEMS04, and Daymet. GZ-METRO records 15-min ridership based on smart card tap-in data at 159 metro stations for 3 months in Guangzhou, China. PEMS04 is a highway traffic flow datasets used in (Song et al. 2020). The dataset records 5-min aggregated traffic flow from the Cal-trans Performance Measurement System (PeMS) (Chen et al. 2001). Daymet is a dataset that provides the daily maximum temperature recordings for Continental North America from 2010 to 2021 (Thornton et al. 2022). The summary of datasets can be found in Table 2. For GZ-METRO and Daymet, 70% of the data were used for training, 20% for testing, and the remaining 10% for validation. For PEMS04, the split ratio is 6:2:2 as used in the original paper (Song et al. 2020). All datasets were applied z-score normalization with statistics obtained from the training set. We selected several spatiotemporal forecasting frameworks as the base models:
  • FC-LSTM (Sutskever, Vinyals, and Le 2014): A Sequence-to-Sequence LSTM model with fully-connected LSTM layers in both encoder and decoder.
  • DCRNN (Li et al. 2018): Diffusion convolution recurrent neural network, which wraps diffusion convolution operation into recurrent neural networks to achieve spatiotemporal forecasting.
  • Graph WaveNet (Wu et al. 2019): A spatiotemporal forecasting model that combines dilated 1D convolution for modeling temporal dynamics and graph convolution for modeling spatial dynamics.
  • AGCRN (Bai et al. 2020): Similar to Graph WaveNet which uses an adaptive adjacency matrix, the model uses GRU to model temporal dependency.
@@
Qualitative Study
@@
(c) ACF without time embedding.
(d) ACF with our method.
Figure 4: (a)–(b) Residual correlation matrices of GZ-METRO at lags $S=0,72,504$, without and with our DMD-based embedding. (c)–(d) Autocorrelation functions (ACF) of 12-step prediction residuals from three sensors, showing how our embedding reduces periodic error peaks.
Effectiveness of DMD-based time embedding
To evaluate the effectiveness of our DMD-based time embedding, we compare it with several alternatives for encoding periodic patterns: (i) time of day (D), (ii) time of day, day of week (DW), and (iii) the learnable sinusoidal embedding method Time2Vec (Kazemi et al. 2019).
Our observations align with the findings of (Tancik et al. 2020), which showed that learned Fourier-based time features tend to concentrate on high-frequency components and struggle to capture long-range periodicity. We further inspected the learned frequencies of Time2Vec and found that they mostly correspond to short-term fluctuations with periods between 5 and 15 time steps (i.e., 25–75 minutes), indicating a failure to model daily or weekly cycles. By contrast, our DMD-based embedding captures multiple frequencies of varying scales, including long-term modes.
Residual correlation and temporal dependency analysis
Figure 4(a) and Figure 4(b) show the residual correlation matrices computed from the 12-step-ahead prediction errors at three time lags: $S=0$, 727272, and 504504504 steps, corresponding to contemporaneous, daily, and weekly intervals. Each panel visualizes the correlations between residual vectors $\bm{\eta}_{t}$ and $\bm{\eta}_{t-S}$ over all spatiotemporal positions. Without our method, we observe strong overall correlations at $S=0$ (0.138), and noticeable dependencies at $S=72$ (0.081) and $S=504$ (0.096). In contrast, our method reduces these values to 0.104, 0.063, and 0.079, respectively. These reductions indicate that our model effectively captures structured periodic patterns and reduces error autocorrelation across periods.
To further validate this observation, we analyze the autocorrelation function (ACF) of the prediction residuals from three representative stations (A, B, and C). Figure 4(c) shows the ACF plots of the original Graph WaveNet model. We observe clear peaks around lag 72 and 144, corresponding to daily and bi-daily periodicity, suggesting that the model fails to fully capture these regular patterns, which manifest as cyclic structure in the residuals. In comparison, Figure 4(d) illustrates the ACF of the residuals using Graph WaveNet augmented with our time embedding. The periodic peaks at lag 72 and 144 are substantially diminished across all three stations. This indicates that the DMD-based time features help the model internalize the dominant periodic components in the data, leading to less temporally correlated residuals and improved forecasting generalization.
Conclusion